OpenStudy (anonymous):

-(5/w+2) = -8 - (7/w-1)

OpenStudy (kenljw):

-(5/w+2) = -8 - (7/w-1) -(5 +2 W) = -8 W - (7-W) -(5 +2 W) = -7 W - 7 -5 +5 W) = - 7 W = - 7/5

OpenStudy (zzr0ck3r):

there should be more than one solution

hero (hero):

\[-\frac{5}{w+2} = -8 -\frac{7}{w-1}\] \[-\frac{5}{w+2} = -\left(8 +\frac{7}{w-1}\right)\] \[\frac{5}{w+2} = \left(8 +\frac{7}{w-1}\right)\] \[\frac{5}{w+2} = \frac{8(w-1)+7}{w-1}\] \[\frac{5}{w+2} = \frac{8w-8+7}{w-1}\] \[\frac{5}{w+2} = \frac{8w-1}{w-1}\] \[5(w-1) = (8w - 1)(w+2)\] \[\frac{w-1}{w+2}=\frac{8w-1}{5}\] \[\frac{w+2-3}{w+2}=\frac{8w-1}{5}\] \[\frac{w+2}{w+2}- \frac{3}{w+2}=\frac{8w-1}{5}\] \[1- \frac{3}{w+2}=\frac{8w-1}{5}\] \[1- \frac{3}{w+2}=\frac{5 + 8w-6}{5}\] \[1- \frac{3}{w+2}=\frac{5}{5} + \frac{8w-6}{5}\] \[1- \frac{3}{w+2}=1 + \frac{8w-6}{5}\] \[-\frac{3}{w+2}=\frac{8w-6}{5}\] \[-(3)(5) = (8w-6)(w+2)\] \[-15 = 8w^2 + 16w - 6w - 12\] \[-15 = 8w^2 + 10w - 12\] \[0 = 8w^2 + 10w + 3\] \[0 = 8w^2 + 6w + 4w + 3\] \[0 = 2w(4w + 3) + 1(4w + 3)\] \[0=(4w+3)(2w + 1)\] Finish the rest

OpenStudy (zzr0ck3r):

the man

OpenStudy (anonymous):

Is it -(5/w+2) = -8 - (7/w-1) or -[5/(w+2)] = -8 - [7/(w-1)] because they are different answers.

hero (hero):

The latter

OpenStudy (anonymous):

If it is as written then it should go like this... -(5/w + 2) = -8 -(7/w - 1) -5/w - 2 = -8 + (-7/w) +1 -5/w - 2 = -7 + (-7/w) -5/w + 5 = -7/w -5 + 5w = -7 5w = -2 w = -2/5 (or 0.4)

hero (hero):

I'm pretty confident that is as I posted. I guess we'll never know.

OpenStudy (anonymous):

Well, either way now they have both.